### v2

# Debiased Nonparametric Regression for Statistical Inference and Distributionally Robustness

Abstract
This study proposes a debiasing method for smooth nonparametric estimators. While machine learning techniques such as random forests and neural networks have demonstrated strong predictive performance, their theoretical properties remain relatively underexplored. Specifically, many modern algorithms lack assurances of pointwise asymptotic normality and uniform convergence, which are critical for statistical inference and robustness under covariate shift and have been well-established for classical methods like Nadaraya-Watson regression. To address this, we introduce a model-free debiasing method that guarantees these properties for smooth estimators derived from any nonparametric regression approach. By adding a correction term that estimates the conditional expected residual of the original estimator, or equivalently, its estimation error, we obtain a debiased estimator with proven pointwise asymptotic normality, and uniform convergence. These properties enable statistical inference and enhance robustness to covariate shift, making the method broadly applicable to a wide range of nonparametric regression problems.
1 Introduction
This paper addresses the problem of nonparametric regression. Let $X\in{\mathcal{X}}$ be a covariate and $Y\in{\mathcal{Y}}$ a target variable, where ${\mathcal{X}}\in[0,1]^{d}$ and ${\mathcal{Y}}\subset{\mathbb{R}}$ represent the covariate and target spaces, respectively. While we primarily focus on the case where $X$ is one-dimensional for simplicity, the results presented in this paper extend to $d$-dimensional covariates.
Let $P$ denote the joint distribution of $(X,Y)$, and define the regression function under $P$ as
$f_{P}(X)={\mathbb{E}}_{P}\big{[}Y\mid X\big{]},$
where ${\mathbb{E}}_{P}\big{[}\cdot\big{]}$ represents the expectation operator with respect to $P$. We assume access to a dataset $\Big{\{}(X_{i},Y_{i})\Big{\}}_{i=1}^{n}$, where $(X_{i},Y_{i})$ are independent and identically distributed samples from a true distribution $P_{0}$. The true regression function is denoted by $f_{0}=f_{P_{0}}$. Our objective is to estimate $f_{0}$ based on the observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$.
Notation.
We denote the distribution induced by a regression function $f$ as $P_{f}$, and the expectation operator under $P_{f}$ as ${\mathbb{E}}_{f}$. For a real-valued vector $z$, let $\|z\|$ denote its Euclidean norm. For a measurable function $h\colon{\mathcal{X}}\to{\mathbb{R}}$, let $\|h\|_{\infty}\coloneqq\sup_{x\in{\mathcal{X}}}|h(x)|$ denote the sup-norm, and $\|h\|_{2}\coloneqq\sqrt{{\mathbb{E}}\left[h(X)^{2}\right]}$ denote the $L^{2}$ norm. t
1.1 Content of this study
In Section 2, we define our debiased estimator in nonparametric regression. Our proposed estimator consists of the following three steps: estimate the regression function $f_{0}$ using any smooth nonparametric estimator; estimate the estimation error of the first stage estimator using the local polynomial regression; combine the first and second stage estimators. Here, we use the smoothness in the meaning that the either of the target nonparametric estimator, the true regression function, or their difference belongs to the Hölder class.
In Section 3, we show that our debiased nonparametric regression estimators have asymptotic normality and uniform convergence under mild conditions, like the smoothness. Furthermore, our estimator has the doubly robust property; that is, either of the first-stage nonparametric regression estimator or the second-stage conditional expected residual estimator is consistent, the resulting debiased estimator is consistent.
1.2 Related work
Doubly robust and debiased estimation has garnered significant attention across various fields, including statistics, economics, epidemiology, and machine learning. Doubly robust estimators are widely used in the context of semiparametric statistics and causal inference (van der Vaart, 1998; Bang & Robins, 2005). With the development of machine learning regression models, debiasing methods utilizing doubly robust estimators have also been extensively studied (Chernozhukov et al., 2018), particularly in conjunction with sample splitting techniques (Klaassen, 1987; van der Vaart, 2002; Zheng & van der Laan, 2011). While much of the existing literature focuses on parametric or semiparametric models, this study extends these approaches to nonparametric regression.
Nonparametric regression in machine learning faces several challenges:
  • For many machine learning estimators, the convergence rate to the true function is unknown, particularly with respect to minimax optimality.
  • Asymptotic normality remains unproven for many machine learning estimators, including random forests (Breiman, 2001).
  • Uniform convergence is not established for many estimators, such as neural networks (Schmidt-Hieber & Zamolodtchikov, 2024).
  • Even for machine learning methods that theoretically satisfy these properties, verifying the required conditions can be challenging.
Minimax optimality and asymptotic normality are crucial in statistical analysis, motivating efforts to establish these properties. For example, Wager & Athey (2018) and Mourtada et al. (2020) demonstrate asymptotic normality and minimax optimality for their modified random forests, respectively, but not for the original random forest proposed by Breiman (2001). These simplifications facilitate desirable theoretical results but may degrade empirical performance.
Uniform convergence is particularly important for robustness under distributional shifts, such as covariate shift (Shimodaira, 2000), where nonparametric regression models are known to be effective. However, uniform convergence is rarely established for modern nonparametric regression estimators, including classical methods like local linear and series regression. This limitation stems from their data-adaptive nature and the prevalent use of empirical process arguments, which focus on population risk measures, such as mean squared error, rather than sup-norm bounds for uniform convergence. Schmidt-Hieber & Zamolodtchikov (2024) addresses this issue by demonstrating restricted uniform optimality for neural network regression in the one-dimensional covariate case, but their approach is not easily generalizable.
Our method shares some motivations with debiasing approaches in high-dimensional regression problems. For instance, the Lasso estimator, while widely used, introduces a bias that diminishes with sample size at a slower rate than $\sqrt{n}$, hindering asymptotic normality (Tibshirani, 1996; Bühlmann & van de Geer, 2011). To mitigate this, approaches such as van de Geer et al. (2014) incorporate bias-correction terms, yielding improved asymptotic properties like normality and efficiency (Janková & van de Geer, 2018).
This study generalizes our earlier work on doubly robust methods for nonparametric regression discontinuity design (Kato, 2024). Independently, Chernozhukov et al. (2024) develops conditional influence functions with similar motivations, building on work by Ichimura & Newey (2022).
From a technical perspective, this study is heavily influenced by Belloni et al. (2011) and Kennedy et al. (2024). Belloni et al. (2011) investigates various properties of nonparametric series estimators, which were subsequently utilized in Kennedy et al. (2024) to study minimax nonparametric estimators for conditional average treatment effects. We leverage these technical findings, alongside classical results from Stone (1980, 1982) and Tsybakov (2008).
2 The debiased estimator
We randomly split the observations ${\mathcal{D}}$ into two datasets, ${\mathcal{D}}^{(1)}$ and ${\mathcal{D}}^{(2)}$, satisfying ${\mathcal{D}}={\mathcal{D}}^{(1)}\cup{\mathcal{D}}^{(2)}$. For simplicity, let $n$ be an even value, and define $m=n/2$. Then, for each $\ell\in\{1,2\}$ we denote
${\mathcal{D}}^{(\ell)}\coloneqq\big{\{}(X_{i},Y_{i})\big{\}}_{i\in{\mathcal{I}}^{(\ell)}},$
where ${\mathcal{I}}^{(\ell)}$ is the index set of ${\mathcal{D}}^{(\ell)}$; that is, ${\mathcal{I}}^{(1)}\cup{\mathcal{I}}^{(2)}=\{1,2,\dots,n\}$.
We consider the following three-stage estimation: for each point $x_{0}\in{\mathcal{X}}$ of interest:
First-stage:
Estimate $f_{0}$ using any smooth model and denote the estimator by $\widehat{f}_{n}$.
Second-stage:
Estimate the conditional expected residual of the first-stage estimator ${\mathbb{E}}\big{[}Y-f(X)\mid X=x_{0}\big{]}$ or, equivalently, the estimation error $f_{0}(x_{0})-\widehat{f}_{n}(x_{0})$. We denote the estimator by
Third-stage:
Summing the first and second stage estimators and obtain
$\widetilde{f}_{n}(x_{0})\coloneqq\widehat{b}_{n}(x_{0})+\widehat{f}_{n}(x_{0}),$ (1)
as a debiased estimator of $f_{0}$.
As explained in Section 3, we can show the asymptotic normality and uniform convergence for the debiased estimator. In contrast, many of the modern machine learning methods lack those properties, even though their performances are guaranteed for the population risk.
In this section, we briefly explain the construction of the first-stage estimator $\widehat{f}_{n}$ and the second-stage estimator $\widehat{b}_{n}$.
2.1 First-stage nonparametric regression
For the first-stage estimator $\widehat{f}_{n}$, we can use any regression methods if $f_{0}-\widehat{f}_{n}$ is smooth. There can be several definitions for the smoothness. In this study, for simplicity, we focus on the smoothness in the meaning of the Hölder class (Definition 3.1). We note that both $f_{0}$ and $\widehat{f}_{n}$ do not have to be smooth if the difference $f_{0}-\widehat{f}_{n}$ is smooth.
We emphasize that in our analysis, we do not require specific properties for $\widehat{f}_{n}$ except for the smoothness. The first-stage estimator $\widehat{f}_{n}$ can converge to the true function $f_{0}$ with a very slow rate or even be inconsistent. Unless the smoothness conditions are satisfied, our asymptotic theoretical results hold, although the empirical finite-sample performance depends on the choice of the first stage estimator.
2.2 Second-stage conditional expected residual estimation
For each element $X_{j}$ of covariates $X=(X_{1}\ X_{2}\ \cdots\ X_{d})^{\top}\in{\mathcal{X}}$, let us define the Legendre polynomial series basis $\widetilde{\rho}\colon{\mathbb{R}}\to{\mathbb{R}}$ as
$\widetilde{\rho}(X_{j})\coloneqq\Big{(}\rho_{0}(X_{j})\ \ \rho_{1}(X_{j})\ \ \cdots\ \ \rho_{\lfloor\ell\rfloor}(X_{j})\Big{)}^{\top},$
where
$\widetilde{\rho}_{m}(X_{j})\coloneqq\sum^{m}_{k=0}(-1)^{\ell+m}\sqrt{2m+1}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}m+k\\ k\end{pmatrix}X^{k}_{j}.$
Then, we define $\rho(X_{i})$ to be the corresponding tensor product of all interactions of $\widetilde{\rho}_{m}(X_{1,i}),\dots,\widetilde{\rho}_{m}(X_{d,i})$.
By using the Legendre polynomial basis, we define the local linear regression as follows:
$\widehat{\beta}_{n}(x)\coloneqq\begin{pmatrix}\widehat{\beta}_{1,n}(x)&\widehat{\beta}_{2,n}(x)&\cdots&\widehat{\beta}_{\ell,n}(x)\end{pmatrix}^{\top}$
$\coloneqq\operatorname*{arg\,min}_{\beta\in{\mathbb{R}}^{\ell+1}}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})-\beta^{\top}\rho\left(\frac{X_{i}-x}{h_{n}}\right)\right)^{2}K\left(\frac{X_{i}-x}{h_{n}}\right),$
where $K\colon{\mathcal{X}}\to{\mathbb{R}}$ is a kernel function, defined as
$K(x)\coloneqq K_{h}(x)\coloneqq\mathbbm{1}\left[\big{\|}x-x_{0}\big{\|}\leq h\right].$
Then, the conditional expected residual estimator corresponds to $\widehat{\beta}_{1,n}(x)$, the first element of $\widehat{\beta}_{n}(x)$; that is,
$\widehat{b}_{n}(x)=\widehat{\beta}_{1,n}(x).$
This second-stage estimator corresponds to the bias-correction term in the construction of semiparametric efficient estimators (Schuler & van der Laan, 2024). Such an estimator is referred to as the bias-correction or one-step estimator in the context of semiparametric statistics (van der Vaart, 2002).
2.3 Discussion
Our estimator (1) is closely related to the influence function in conditional mean estimation (Kennedy, 2023). This type of estimator is also connected to the Neyman orthogonal score (Chernozhukov et al., 2018), which is almost mathematically equivalent to the canonical gradient in the one-step bias corection (Schuler & van der Laan, 2024). Note that in conditional mean estimation, the construction of the influence function becomes complicated due to the conditioning. For the influence function including the parameter represented via the conditional expected value, see Ichimura & Newey (2022). Independently and simultaneously of us, Chernozhukov et al. (2024) investigates a related topic, with different interest and method.
3 Convergence analysis
This section presents convergence analysis of $\widetilde{f}_{n}$. First, we define the Hölder class as follows:
Definition 3.1 (Hölder class).
Let ${\mathcal{T}}$ be an interval in ${\mathbb{R}}^{d}$ and let $s$ and $L$ be two positive numbers. Let $D^{\alpha}=\frac{\partial^{\alpha}}{\partial x_{1}^{\alpha_{1}}\dots\partial x_{d}^{\alpha_{d}}}$ be the partial derivative operator. The Hölder class $\Sigma(s,L)$ on ${\mathcal{T}}$ is defined as the set of $\ell=\lfloor s\rfloor$ times differentiable function $f\colon{\mathcal{T}}\to{\mathbb{R}}$ satisfying the following conditions:
  • Its derivatives up to order $\lfloor\ell\rfloor$ are bounded as
    $\left|D^{\alpha}f(x)\right|\leq C<\infty\qquad\forall x\in\mathcal{X},$
    for all $\alpha=(\alpha_{1},\dots,\alpha_{d})$ with $\sum_{j}\alpha_{j}\leq\ell$.
  • Its $\ell$-order derivatives satisfying the Lipschitz condition: there exists a constant $0<C\infty$ such that
    $\left|D^{\beta}f(x)-D^{\beta}f(x^{\prime})\right|\leq C\|x-x^{\prime}\|^{s-\ell}\qquad\forall x,x^{\prime}\in\mathcal{X}$
    for all $\beta=(\beta_{1},\dots,\beta_{d})$ with $\sum_{j}\beta_{j}=\ell$.
Define
^$\widehat{{\mathcal{B}}}_{n}(x_{0})$ $\coloneqq\frac{1}{mh^{d}_{n}}\sum_{i\in{\mathcal{I}}^{(2)}}\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{X_{i}-x_{0}}{h_{n}}\right),$
$\widehat{w}_{h}(x,x_{0})$ $\coloneqq\frac{1}{h^{d}_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(\frac{x-x_{0}}{h_{n}}\right)K\left(\frac{x-x_{0}}{h_{n}}\right).$
Then, the second-stage estimator is given as
$\widehat{b}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0}).$
3.1 Bias and variance decomposition
First, we investigate the bias and variance of $\widetilde{f}_{n}$. That is, we aim to analyze the bias and variance of $\widetilde{b}_{n}(x_{0})$, denoted as
$\mathrm{Bias}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\right]-f_{0}(x_{0}),$
$\mathrm{Variance}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\right]\Big{)}^{2}\right].$
Let us introduce the population values of ^$\widehat{{\mathcal{B}}}_{n,x}$ and $\widehat{s}_{n,x}$ as
~$\widetilde{{\mathcal{B}}}_{n}(x_{0})$ $\coloneqq\int\rho\left(\frac{x-x_{0}}{h_{n}}\right)\rho\left(\frac{x-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{x-x_{0}}{h_{n}}\right){\mathrm{d}}F(x).$
For this value, the covariate distribution, and the smoothness of $f_{0}$ and $\widehat{f}_{n}$, we make the following assumptions.
Assumption 3.2.
Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For a point $x_{0}\in{\mathcal{X}}$ of interest, the following holds:
  • $f_{0}-\widehat{f}$ belongs to the Hölder class $\Sigma(s,L)$.
  • The eigenvalues of ^$\widehat{{\mathcal{B}}}_{n}(x)$ is bounded above and below away from zero.
  • There exists a constant $C>0$ such that $\int\mathbbm{1}\big{[}\|x-x_{0}\|\leq h\big{]}{\mathrm{d}}P(x)\leq Ch^{d}$,
where $P(x)$ denotes a distribution of $x$.
Theorem 3.3 (Bias and variance decomposition).
Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. For the values, suppose that Assumptions 3.2 holds. Let $h_{n}$ be the bandwidth of the local polynomial estimator. Then, for any $\varepsilon>0$ and for all $x_{0}\in[0,1]$, there exits $n_{0}>0$ such that for all $n\geq n_{0}$, with probability $1-\varepsilon$, it holds that
$\left|\mathrm{Bias}(x_{0})\right|\leq C_{1}h^{s}_{n},$
$\mathrm{Variance}(x_{0})\leq\frac{C_{2}}{nh^{d}_{n}}+C_{3}h^{2s}_{n},$
where $C_{1},C_{2},C_{3}>0$ are constants independent of $f_{0}$ and $n$.
The smoothness condition on $f_{0}-\widehat{f}$ can be satisfied in several ways. We discuss this point in Section 4.
3.2 Pointwise convergence of the mean squared errors
Theorem 3.3 about the convergences of bias and variance terms directly indicate the convergence of the mean squared errors. This is because we have
${\mathbb{E}}\left[\left(\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=\mathrm{Bias}(x_{0})^{2}+\mathrm{Variance}(x_{0})$
To minimize the mean squared error, we set the bandwidth as
$h_{n}=\alpha n^{-\frac{1}{2s+d}},$
where $\alpha$ is a constant independent of $f_{0}$ and $n$, and recall that $s$ is the parameter of smoothness and $d$ is the dimension of $X$. Then, we state the pointwise convergence of the mean squared errors as follows:
Theorem 3.4 (Pointwise mean squared error convergence).
Suppose that the conditions in Theorem 3.3 holds. Let $h_{n}=\alpha n^{-\frac{1}{2s+d}}$. Then, for all $x_{0}\in[0,1]$, it holds that
$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}\sup_{x_{0}\in{\mathcal{X}}}{\mathbb{E}}_{f}\left[\psi^{-2}_{n}\Big{|}\widetilde{f}_{n}(x_{0})-f(x_{0})\Big{|}^{2}\right]\leq C<\infty,$
where $\psi_{n}\coloneqq n^{-\frac{s}{2s+d}}$ is the rate of convergence and $C$ is a constant depending only on $s$, $L$, $\lambda_{0}$, $a_{0}$, $\sigma^{2}_{\max}$, $K_{\max}$, and $s$.
Theorem 3.4 can be extend to the following mean squared error convergence over the distribution of $X$.
Corollary 3.5 (Mean squared errors over the distribution of $X$).
Suppose that the conditions in Theorem 3.4 holds. Then, it holds that
$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}{\mathbb{E}}_{f}\left[\psi^{-2}_{n}\Big{\|}\widetilde{f}_{n}-f\Big{\|}^{2}_{2}\right]\leq C<\infty,$
where $\psi_{n}\coloneqq n^{-\frac{s}{2s+d}}$ is the rate of convergence and $C$ is a constant depending only on $s$, $L$, $\lambda_{0}$, $a_{0}$, $\sigma^{2}_{\max}$, $K_{\max}$, and $s$.
3.3 Pointwise asymptotic normality
We now establish the asymptotic normality of $\widetilde{f}(x_{0})$. This result follows directly from Theorem 3.3.
Theorem 3.6 (Asymptotic normality).
Suppose one of the upper bounds in Theorem 3.4 holds. Then, it tholds that
$\sqrt{nh^{d}_{n}}\Big{(}\widetilde{f}(x_{0})-f_{0}(x_{0})\Big{)}\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V(x_{0})),$
where $V(x_{0})\coloneqq nh^{d}_{n}\mathrm{Variance}(x_{0})\leq C_{2}+o(h_{n})$.
The asymptotic normality result facilitates statistical inference. However, constructing a confidence interval is challenging due to the unknown parameter $s$ involved in $h_{n}$. To address this issue, the concept of honest confidence intervals has been developed. For example, the bootstrap method can be employed to obtain valid confidence intervals, even when certain parameters remain unspecified.
3.4 Uniform convergence
Lastly, we establish the uniform (or sup-norm) convergence of $\widetilde{f}(x_{0})$. Uniform convergence bounds the supremum of the estimation error uniformly over $X\in{\mathcal{X}}$, without relying on the distribution of $X$, whereas mean squared convergence depends on the distribution of $X$. This theoretical guarantee is particularly useful for addressing distribution shift problems, such as covariate shift (Shimodaira, 2000). In covariate shift scenarios, the distribution of $X$ differs between the training and test data. While mean squared convergence may be affected by such a shift, uniform convergence remains unaffected, ensuring robust performance for both training and test data. Consequently, uniform convergence is critical in discussions involving distribution shifts.
Theorem 3.7 (Uniform convergence).
Suppose that $f_{0}$ belongs to a Hölder class $\Sigma(s,L)$ on $\mathcal{X}$, where $s>0$ and $L>0$. Let $\widehat{f}_{n}$ be the $\ell$) estimator of order $\ell=\lfloor s\rfloor$ with bandwidth
$h_{n}\coloneqq\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+d}}$
for some $\alpha>0$. Suppose the following:
  • Assumption 3.2 holds.
  • $\varepsilon_{i}$ is a sub-Gaussian random variable satisfying
    ${\mathbb{E}}\big{[}\exp\big{(}\lambda\varepsilon_{i}\big{)}\big{]}\leq\exp\big{(}K^{2}\lambda^{2}\big{)},\quad\forall\lambda\in{\mathbb{R}},$
    for some constant $K>0$.
  • $K$ is a Lipschitz kernel: $K\in\Sigma(1,L_{K})$ on $\cal{X}$ with $0<L_{K}<\infty$.
Then, there exists a constant $C<\infty$ such that
$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}{\mathbb{E}}_{f}\big{[}\psi^{-2}_{n}\big{\|}\widetilde{f}_{n}-f\big{\|}^{2}_{\infty}\big{]}\leq C,$
where
$\psi_{n}\coloneqq\left(\frac{\log(n)}{n}\right)^{\frac{s}{2s+d}}.$
Uniform convergence provides robustness against distributional shifts but has not been proven for many nonparametric regression estimators produced by machine learning methods. This limitation arises partly from their data-adaptive nature and the reliance on empirical process techniques within the empirical risk minimization framework (van de Geer, 2009). Using empirical processes and model complexity, estimation error can be evaluated in a general manner, avoiding significant dependence on specific algorithms. However, such approaches often focus on population risk, such as mean squared error, without addressing how these results translate into uniform convergence guarantees.
For example, Schmidt-Hieber & Zamolodtchikov (2024) tackle this issue by proposing an estimator based on neural networks, demonstrating restricted uniform convergence in the case of one-dimensional covariates. In contrast, our proposed method is model-free and accommodates multiple-dimensional covariates, offering broader applicability and robustness.
3.5 Double robustness
Our proposed debiased estimator possesses the property of double robustness; that is, if either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, the resulting debiased estimator $\widetilde{f}_{n}$ is also consistent.
Recall that the debiased estimator is defined as
$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$
Suppose that $\widehat{f}_{n}(x_{0})$ converges in probability to $f^{\dagger}(x_{0})$ and $\widehat{b}_{n}(x_{0})$ converges in probability to $b^{\dagger}(x_{0})$. We examine the behavior of $\widetilde{f}_{n}(x_{0})$ under two cases: $f^{\dagger}(x_{0})=f_{0}(x_{0})$ and $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$.
First, consider the case where $f^{\dagger}(x_{0})=f_{0}(x_{0})$ and $b^{\dagger}(x_{0})\neq f_{0}(x_{0})-f^{\dagger}(x_{0})$. Then, it follows that
$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-f_{0}(X_{i})\Big{)}w^{\dagger}(X_{i},x_{0})+f_{0}(x_{0})+o_{p}(1),$
where $w^{\dagger}(X_{i},x_{0})$ is a function depending only on $X_{i}$ and $x_{0}$. Since the first term, $\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-f_{0}(X_{i})\Big{)}w^{\dagger}(X_{i},x_{0})$, converges in probability to zero, $\widetilde{f}_{n}(x_{0})$ converges in probability to $f_{0}(x_{0})$.
Next, consider the case where $f^{\dagger}(x_{0})\neq f_{0}(x_{0})$ and $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$. Then, it holds that
$\widetilde{f}_{n}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})+f^{\dagger}(x_{0})+o_{p}(1)=f_{0}(x_{0})+o_{p}(1).$
Thus, in either case, $\widetilde{f}_{n}(x_{0})$ converges in probability to $f_{0}(x_{0})$. We summarize this property in the following theorem.
Theorem 3.8 (Double robustness).
If either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, then it holds that
$\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})\qquad\mathrm{as}\ \ n\to\infty.$
4 Smoothness is all you need
As demonstrated above, our method requires only the smoothness of $\widehat{f}-f_{0}$. This assumption is satisfied when $f_{0}$ and $\widehat{f}$ are individually smooth, but it also holds in cases where $f_{0}$ and $\widehat{f}$ are not smooth, provided that their difference is smooth. Neural networks belong to this class under the analysis of (Schmidt-Hieber, 2020).
Remarkably, our results require only the smoothness of the first-stage estimator (the difference between $\widehat{f}_{n}$ and $f_{0}$). This assumption significantly broadens the applicability of our method. Even when $f_{0}$ and $\widehat{f}_{n}$ are not smooth, our method can still guarantee mean squared convergence, asymptotic normality, and uniform convergence. For instance, Imaizumi & Fukumizu (2019) discusses the convergence of nonparametric estimators using neural networks when $f_{0}$ belongs to a piecewise Hölder class. Under such a case, the liner estimators, including locally polynomial regression, are suboptimal. However, even under the case, the difference $\widehat{f}-f_{0}$ can be smooth, and our smoothness condition still holds.
5 Conclusion
This study developed a debiased estimator for nonparametric regression. For any smooth regression estimator, our debiased estimator gives the following desirable property: mean-squared convergence, asymptotic normality, uniform convergence, and double robustness. The asymptotic normality is useful for statistical inference, while the uniform convergence makes the nonparametric regression estimator distributionally robust. Additionally, either the first-stage or the second-stage nonparametric estimator is consistent, the resulting regression estimator is consistent; that is, doubly robust. Surprisingly, our estimator only requires the smoothness of the first-stage regression estimator. More specifically, if either (i) both the first-stage estimator $\widehat{f}$ and the true regression function $f_{0}$ are smooth or (ii) the difference $\widehat{f}-f_{0}$ is smooth, we can show the above desirable properties for the debiased estimator.
6 Proofs
This section provides the proofs in Section 3.
6.1 Proof of Theorem 3.6
Define
$\mathrm{Bias}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)$ $\coloneqq{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]-f_{0}(x_{0}),$
$\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)$ $\coloneqq{\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{)}^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right],$
$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$ $\coloneqq{\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}.$
Here, the followings hold:
$\mathrm{Bias}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Bias}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right],$
$\mathrm{Variance}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right]+\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right).$
We aim to show that
$\mathrm{Bias}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\leq\frac{L}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(X_{i},x_{0})\Big{|},$ (2)
$\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)=\frac{C_{2}}{nh^{d}_{n}}+o(1),$ (3)
$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)=2C^{2}_{1}h^{2s}.$ (4)
If (2) holds, then we have
${\mathbb{E}}\left[\frac{L}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(X,x_{0})\Big{|}\right]$ $=\int L\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(x,x_{0})\Big{|}{\mathrm{d}}P(x){\mathrm{d}}x$
$=\int L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u,$
where $\frac{x-x_{0}}{h}=u$. We have
$\int L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u$
$=L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\int\left|\frac{1}{h^{d}_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(u\right)\mathbbm{1}\left[\|u\|\leq 1\right]\right|{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u.$
6.1.1 Preliminaries
As preliminary, we first state the following lemma about the properties of the local polynomial estimator, as well as Proposition 1.12 in Tsybakov (2008) and Appendix B.3 in Kennedy et al. (2024).
Lemma 6.1 (From Proposition 1.12 in Tsybakov (2008)).
Let $x_{0}$ be a real number such that ${\mathcal{B}}_{n}(x_{0})>0$ and let $Q$ be a polynomial whose degree is less than or equal to $\ell$. Then, the $\ell$) weights $\widehat{w}_{n}(X_{i},x_{0})$ satisfy
$\sum_{i\in{\mathcal{I}}^{(2)}}Q(X_{i})\widehat{w}_{h}(X_{i},x_{0})=Q(x_{0}).$
for any sample $(X_{1},\dots,X_{n})$.
Lemma 6.2.
Under Assumption 3.2, for any $\varepsilon$, there exists $n_{0}$ such that for all $n\geq n_{0}$, $h\geq\frac{1}{2n}$, and $x_{0}\in\mathcal{X}$, the weights $\widehat{w}_{n}(x,x_{0})$ of the $\ell$) estimator satisfy the followings:
  • $\sup_{x,x_{0}}\Big{|}\widehat{w}_{h}(x,x_{0})\Big{|}\leq\frac{C_{*}}{nh}$ with probability $1-\varepsilon$;
  • ${\mathbb{E}}\left[\sum_{i\in{\mathcal{I}}^{(2)}}\Big{|}\widehat{w}_{n}(X_{i},x_{0})\Big{|}\right]\leq C_{*}+\varepsilon$;
  • $\widehat{w}_{h}(x,x_{0})=0$ if $|x-x_{0}|>h$,
where the constant $C_{*}$ depends only on $\lambda_{0}$, $a_{0}$, and $K_{\max}$.
6.1.2 Bounding the bias (proof of (2))
Proof.
We decompose the bias term as
$\mathrm{Bias}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})$
$={\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]-f_{0}(x_{0})$
$={\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]-f_{0}(x_{0}).$
We have
${\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$={\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}f_{0}(X_{0})-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right].$
Let $\widehat{g}_{n}\coloneqq\widehat{f}_{n}-f_{0}$. Since $\widehat{g}_{n}$ belongs to the Hölder class $\Sigma(s,L)$, we have
${\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}\widehat{g}_{n}(X_{i})-\widehat{g}_{n}(x_{0})\Big{)}\widehat{w}_{h}(X_{i},x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$={\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{\widehat{g}^{(\ell)}(x_{0}+\tau_{i}(X_{i}-x_{0}))-\widehat{g}^{(\ell)}(x_{0})}{\ell!}\big{(}X_{i}-x_{0}\big{)}^{\ell}\widehat{w}_{h}(X_{i},x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right].$
where $0\leq\tau_{i}\leq 1$. Here, we used the Taylor expansion of $\widehat{g}$, and Lemma 6.1.
With probability one, we have
${\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{\widehat{g}^{(\ell)}(x_{0}+\tau_{i}(X_{i}-x_{0}))-\widehat{g}^{(\ell)}(x_{0})}{\ell!}\big{(}X_{i}-x_{0}\big{)}^{\ell}\widehat{w}_{h}(X_{i},x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$\leq\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{L\big{\|}X_{i}-x_{0}\big{\|}^{s}}{\ell!}\Big{|}\widehat{w}_{h}(X_{i},x_{0})\Big{|}$
$=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{L\big{\|}X_{i}-x_{0}\big{\|}^{s}}{\ell!}\Big{|}\widehat{w}_{h}(X_{i},x_{0})\Big{|}\mathbbm{1}\big{[}|X_{i}-x_{0}|\leq h\big{]}$
$\leq\frac{L}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(X_{i},x_{0})\Big{|}.$
6.1.3 Bounding the variance: part I (proof of (3))
Proof.
We have
$\mathrm{Variance}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})$
$={\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-f_{0}(x_{0})\Big{)}^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$={\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{)}^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$={\mathbb{E}}\Bigg{[}\Bigg{(}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(X_{i})$
$\ \ \ \ \ \ \ \ \ \ \ \ -{\mathbb{E}}\left[\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(X_{i})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Bigg{)}^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Bigg{]}$
$={\mathbb{E}}\left[\left(\sum_{i\in{\mathcal{I}}^{(2)}}\xi_{i}\widehat{w}_{h}(X_{i},x_{0})\right)^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$=\sum_{i\in{\mathcal{I}}^{(2)}}\left(\widehat{w}_{h}(X_{i},x_{0})\right)^{2}{\mathbb{E}}\big{[}\xi^{2}_{i}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\big{]}.$
From Assumption 3.2, we have
${\mathbb{E}}\left[\mathrm{Variance}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})\right]$
$\leq\frac{1}{m^{2}}{\mathbb{E}}\left[\sum_{i\in{\mathcal{I}}^{(2)}}\left(\widehat{w}_{h}(X,x_{0})\right)^{2}{\mathbb{E}}\big{[}\xi^{2}_{i}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\big{]}\right]$
$\leq\sigma^{2}_{\max}\frac{1}{m}{\mathbb{E}}\left[\sup_{x,x_{0}}\Big{|}\widehat{w}_{h}(x,x_{0})\Big{|}\right]{\mathbb{E}}\left[\Big{|}\widehat{w}_{h}(X,x_{0})\Big{|}\right]$
$\leq\frac{\sigma^{2}_{\max}C^{2}_{*}}{mh^{d}_{n}}+o(1)$
$=\frac{C_{2}}{nh^{d}_{n}}+o(1).$
6.1.4 Bounding the variance: part II (proof of (4))
Proof.
We have
$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
$={\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-f_{0}(x_{0})+f_{0}(x_{0})-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}$
$={\mathbb{E}}\Big{[}\Big{(}\mathrm{Bias}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})-\mathrm{Bias}(x_{0})\Big{)}^{2}\Big{]}$
$\leq 2C^{2}_{1}h^{2s}_{n}.$
6.2 Proof of Theorem 3.7
Proof.
We decompose the sup-norm as follows:
${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-f_{0}\Big{\|}^{2}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$\leq{\mathbb{E}}\left[2\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}^{2}_{\infty}+2\Big{\|}{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]-f_{0}\Big{\|}^{2}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$\leq 2{\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}^{2}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]+2\left(\sup_{x\in{\mathcal{X}}}\Big{|}\mathrm{Bias}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})\Big{|}_{\infty}\right)^{2}$
$\leq 2{\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]+2C^{2}_{1}h^{2s}_{n}.$
Here, note that
${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
$={\mathbb{E}}\left[\sup_{x_{0}\in{\mathcal{X}}}\left|\sum_{i\in{\mathcal{I}}^{(2)}}\xi_{i}\widehat{w}_{h}(X_{i},x_{0})\right|^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right].$
Hereafter, the proof step is the same as the one of Theorem 1.8 of Tsybakov (2008). We use the condition that $\varepsilon_{i}$ is sub-Gaussian with Corollary 1.3 and Lemma 1.6 in Tsybakov (2008). In the original Corollary 1.3 and Lemma 1.6, Tsybakov (2008) assumes that $\varepsilon_{i}$ is Gaussian, but we can generalize the results for sub-Gaussian random variables with minor modifications.
Finally, we have
${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh^{d}_{n}},$
where $C_{3}>0$ is a constant independent of $f$ and $n$.
We have
${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-f_{0}\Big{\|}^{2}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh}+2C^{2}_{1}h^{2s}_{n}.$
By choosing the bandwidth as
$h_{n}=\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+d}},$
we complete the proof. ∎

### v3

@@
 While machine learning techniques such as random forests and neural networks have demonstrated strong predictive performance, their theoretical properties remain relatively underexplored.
-Specifically, many modern algorithms lack assurances of pointwise asymptotic normality and uniform convergence, which are critical for statistical inference and robustness under covariate shift and have been well-established for classical methods like Nadaraya-Watson regression.
-To address this, we introduce a model-free debiasing method that guarantees these properties for smooth estimators derived from any nonparametric regression approach.
-By adding a correction term that estimates the conditional expected residual of the original estimator, or equivalently, its estimation error, we obtain a debiased estimator with proven pointwise asymptotic normality, and uniform convergence.
-These properties enable statistical inference and enhance robustness to covariate shift, making the method broadly applicable to a wide range of nonparametric regression problems.
+In particular, many modern algorithms lack guarantees of pointwise and uniform risk convergence, as well as asymptotic normality.
+These properties are essential for statistical inference and robust estimation and have been well-established for classical methods such as Nadaraya-Watson regression.
+To ensure these properties for various nonparametric regression estimators, we introduce a model-free debiasing method.
+By incorporating a correction term that estimates the conditional expected residual of the original estimator, or equivalently, its estimation error, into the initial nonparametric regression estimator, we obtain a debiased estimator that satisfies pointwise and uniform risk convergence, along with asymptotic normality, under mild smoothness conditions.
+These properties facilitate statistical inference and enhance robustness to covariate shift, making the method broadly applicable to a wide range of nonparametric regression problems.
 1 Introduction
-This paper addresses the problem of nonparametric regression.
-Let $X\in{\mathcal{X}}$ be a covariate and $Y\in{\mathcal{Y}}$ a target variable, where ${\mathcal{X}}\in[0,1]^{d}$ and ${\mathcal{Y}}\subset{\mathbb{R}}$ represent the covariate and target spaces, respectively.
-While we primarily focus on the case where $X$ is one-dimensional for simplicity, the results presented in this paper extend to $d$-dimensional covariates.
+This study investigates the problem of nonparametric regression.
+While modern machine learning methods enable the estimation of complex regression functions, they often lack the theoretical guarantees that traditional estimators possess, such as pointwise and uniform mean squared error (MSE) convergence, as well as asymptotic normality.
+Asymptotic normality is crucial for statistical inference, whereas uniform risk convergence plays an important role in prediction under covariate shift.
+In this study, we propose a debiasing method for regression estimators and show that, under mild smoothness conditions, they achieve asymptotic normality and uniform convergence.
+Here, we formulate the problem.
+Let $X\in{\mathcal{X}}$ be a $d$-dimensional covariate and $Y\in{\mathcal{Y}}$ be a target variable, where ${\mathcal{X}}\subseteq{\mathbb{R}}^{d}$ and ${\mathcal{Y}}\subset{\mathbb{R}}$ represent the covariate and target spaces, respectively.
 Let $P$ denote the joint distribution of $(X,Y)$, and define the regression function under $P$ as
@@
 where ${\mathbb{E}}_{P}\big{[}\cdot\big{]}$ represents the expectation operator with respect to $P$.
-We assume access to a dataset $\Big{\{}(X_{i},Y_{i})\Big{\}}_{i=1}^{n}$, where $(X_{i},Y_{i})$ are independent and identically distributed samples from a true distribution $P_{0}$.
+We assume access to observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$, where $(X_{i},Y_{i})$ are independent and identically distributed samples from a true distribution $P_{0}$.
 The true regression function is denoted by $f_{0}=f_{P_{0}}$.
-Our objective is to estimate $f_{0}$ based on the observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$.
+Our objective is to estimate $f_{0}$ from the observations $\big{\{}(X_{i},Y_{i})\big{\}}_{i=1}^{n}$, aiming for pointwise and uniform risk convergence, as well as asymptotic normality.
 Notation.
-We denote the distribution induced by a regression function $f$ as $P_{f}$, and the expectation operator under $P_{f}$ as ${\mathbb{E}}_{f}$.
 For a real-valued vector $z$, let $\|z\|$ denote its Euclidean norm.
 For a measurable function $h\colon{\mathcal{X}}\to{\mathbb{R}}$, let $\|h\|_{\infty}\coloneqq\sup_{x\in{\mathcal{X}}}|h(x)|$ denote the sup-norm, and $\|h\|_{2}\coloneqq\sqrt{{\mathbb{E}}\left[h(X)^{2}\right]}$ denote the $L^{2}$ norm.
-t
 1.1 Content of this study
-In Section 2, we define our debiased estimator in nonparametric regression.
-Our proposed estimator consists of the following three steps: estimate the regression function $f_{0}$ using any smooth nonparametric estimator; estimate the estimation error of the first stage estimator using the local polynomial regression; combine the first and second stage estimators.
-Here, we use the smoothness in the meaning that the either of the target nonparametric estimator, the true regression function, or their difference belongs to the Hölder class.
-In Section 3, we show that our debiased nonparametric regression estimators have asymptotic normality and uniform convergence under mild conditions, like the smoothness.
-Furthermore, our estimator has the doubly robust property; that is, either of the first-stage nonparametric regression estimator or the second-stage conditional expected residual estimator is consistent, the resulting debiased estimator is consistent.
+In Section 2, we define our debiased estimator for nonparametric regression.
+Our proposed estimator consists of the following three steps: (1) estimate the regression function $f_{0}$ using a nonparametric regression method; (2) estimate the expected conditional residual, or equivalently, the estimation error of the first-stage estimator, using local polynomial regression; and (3) add the second-stage estimator to the first-stage estimator.
+In Section 3, we present an example of our proposed debiased estimator, applying local polynomial regression for conditional expected residual estimation.
+In Section 4, we establish that our debiased nonparametric regression estimator achieves pointwise and uniform MSE convergence if either (a) the first-stage nonparametric regression estimator and the true regression function, or (b) their difference, belongs to the Hölder class.
+Using pointwise MSE convergence, we also prove asymptotic normality.
+Furthermore, our estimator possesses the doubly robust property: if either the first-stage nonparametric regression estimator or the second-stage conditional expected residual estimator is consistent, then the resulting debiased estimator is also consistent.
+The key takeaway is that our debiased estimator ensures desirable theoretical properties, typically associated with classical methods, for a wide range of modern first-stage estimators.
+As long as certain smoothness conditions are met, our method effectively corrects bias in a broad class of initial estimators, equipping the resulting estimator with strong theoretical guarantees.
 1.2 Related work
-Doubly robust and debiased estimation has garnered significant attention across various fields, including statistics, economics, epidemiology, and machine learning.
-Doubly robust estimators are widely used in the context of semiparametric statistics and causal inference (van der Vaart, 1998; Bang & Robins, 2005).
-With the development of machine learning regression models, debiasing methods utilizing doubly robust estimators have also been extensively studied (Chernozhukov et al., 2018), particularly in conjunction with sample splitting techniques (Klaassen, 1987; van der Vaart, 2002; Zheng & van der Laan, 2011).
+Doubly robust and debiased estimation has attracted significant attention across various fields, including statistics (Tsiatis, 2007), economics (Chernozhukov et al., 2018), epidemiology (Bang & Robins, 2005), and machine learning (Kallus & Uehara, 2020).
+With the advancement of machine learning regression models, debiasing methods based on doubly robust estimators have been extensively studied (Chernozhukov et al., 2018), particularly in conjunction with sample splitting techniques (Klaassen, 1987; van der Vaart, 2002; Zheng & van der Laan, 2011).
 While much of the existing literature focuses on parametric or semiparametric models, this study extends these approaches to nonparametric regression.
-Nonparametric regression in machine learning faces several challenges:
-  • For many machine learning estimators, the convergence rate to the true function is unknown, particularly with respect to minimax optimality.
-  • Asymptotic normality remains unproven for many machine learning estimators, including random forests (Breiman, 2001).
-  • Uniform convergence is not established for many estimators, such as neural networks (Schmidt-Hieber & Zamolodtchikov, 2024).
-  • Even for machine learning methods that theoretically satisfy these properties, verifying the required conditions can be challenging.
-Minimax optimality and asymptotic normality are crucial in statistical analysis, motivating efforts to establish these properties.
-For example, Wager & Athey (2018) and Mourtada et al.
-(2020) demonstrate asymptotic normality and minimax optimality for their modified random forests, respectively, but not for the original random forest proposed by Breiman (2001).
-These simplifications facilitate desirable theoretical results but may degrade empirical performance.
-Uniform convergence is particularly important for robustness under distributional shifts, such as covariate shift (Shimodaira, 2000), where nonparametric regression models are known to be effective.
-However, uniform convergence is rarely established for modern nonparametric regression estimators, including classical methods like local linear and series regression.
-This limitation stems from their data-adaptive nature and the prevalent use of empirical process arguments, which focus on population risk measures, such as mean squared error, rather than sup-norm bounds for uniform convergence.
-Schmidt-Hieber & Zamolodtchikov (2024) addresses this issue by demonstrating restricted uniform optimality for neural network regression in the one-dimensional covariate case, but their approach is not easily generalizable.
-Our method shares some motivations with debiasing approaches in high-dimensional regression problems.
-For instance, the Lasso estimator, while widely used, introduces a bias that diminishes with sample size at a slower rate than $\sqrt{n}$, hindering asymptotic normality (Tibshirani, 1996; Bühlmann & van de Geer, 2011).
-To mitigate this, approaches such as van de Geer et al.
-(2014) incorporate bias-correction terms, yielding improved asymptotic properties like normality and efficiency (Janková & van de Geer, 2018).
-This study generalizes our earlier work on doubly robust methods for nonparametric regression discontinuity design (Kato, 2024).
-Independently, Chernozhukov et al.
-(2024) develops conditional influence functions with similar motivations, building on work by Ichimura & Newey (2022).
-From a technical perspective, this study is heavily influenced by Belloni et al.
-(2011) and Kennedy et al.
-(2024).
-Belloni et al.
-(2011) investigates various properties of nonparametric series estimators, which were subsequently utilized in Kennedy et al.
-(2024) to study minimax nonparametric estimators for conditional average treatment effects.
-We leverage these technical findings, alongside classical results from Stone (1980, 1982) and Tsybakov (2008).
+The application of debiased machine learning to nonparametric regression is discussed in Kennedy (2023), though their focus differs from ours.
+Nonparametric regression in machine learning presents several challenges:
+  • Pointwise and uniform MSE convergence is not established for many estimators, such as neural networks and random forests.
+  • Asymptotic normality remains unproven for numerous machine learning estimators.
+Pointwise and uniform MSE convergence and asymptotic normality are critical in statistical analysis, motivating significant efforts to establish these properties.
+For instance, Wager & Athey (2018) and Mourtada et al.
+(2020) demonstrate asymptotic normality and optimality, respectively, for their modified random forests, but not for the original random forest of Breiman (2001).
+Pointwise and uniform convergence is particularly relevant for robustness under distributional shifts, such as covariate shift (Shimodaira, 2000).
+These convergence properties are rarely established for modern nonparametric estimators, whereas classical methods like local linear and series regression possess them.
+This gap arises in part due to the data-adaptive nature of modern methods and the frequent use of empirical process arguments, which often focus on population MSE, where expectations are taken over the covariate distribution rather than on pointwise or sup-norm MSE.
+Schmidt-Hieber & Zamolodtchikov (2024) addresses this concern by demonstrating restricted uniform optimality for neural network regression in the one-dimensional covariate setting, but this approach does not generalize easily.
+Our method shares some motivations with debiasing approaches in high-dimensional regression.
+For example, the Lasso estimator, while widely used, introduces a bias that decreases with sample size at a slower rate than $\sqrt{n}$, thereby hindering asymptotic normality (Tibshirani, 1996; Bühlmann & van de Geer, 2011).
+To mitigate this, van de Geer et al.
+(2014) proposes incorporating bias-correction terms, leading to improved asymptotic properties such as normality and efficiency (Janková & van de Geer, 2018).
+This study generalizes our previous work on doubly robust methods for nonparametric regression discontinuity design (Kato, 2024).
+Concurrently with our work, Chernozhukov et al.
+(2024) and van der Laan et al.
+(2025) also explore double machine learning in the context of nonparametric regression, though their emphases differ substantially from ours.
+For example, Chernozhukov et al.
+(2024) develop conditional influence functions, which are related to the framework of Ichimura & Newey (2022).
 2 The debiased estimator
-We randomly split the observations ${\mathcal{D}}$ into two datasets, ${\mathcal{D}}^{(1)}$ and ${\mathcal{D}}^{(2)}$, satisfying ${\mathcal{D}}={\mathcal{D}}^{(1)}\cup{\mathcal{D}}^{(2)}$.
-For simplicity, let $n$ be an even value, and define $m=n/2$.
-Then, for each $\ell\in\{1,2\}$ we denote
+In this section, we define our debiased estimator in a general way.
+First, we randomly split the observations ${\mathcal{D}}$ into two datasets, ${\mathcal{D}}^{(1)}$ and ${\mathcal{D}}^{(2)}$, such that ${\mathcal{D}}={\mathcal{D}}^{(1)}\cup{\mathcal{D}}^{(2)}$.
+For simplicity, let $n$ be even and set $m=n/2$.
+Then, for each $\ell\in\{1,2\}$, we define
 ${\mathcal{D}}^{(\ell)}\coloneqq\big{\{}(X_{i},Y_{i})\big{\}}_{i\in{\mathcal{I}}^{(\ell)}},$
-where ${\mathcal{I}}^{(\ell)}$ is the index set of ${\mathcal{D}}^{(\ell)}$; that is, ${\mathcal{I}}^{(1)}\cup{\mathcal{I}}^{(2)}=\{1,2,\dots,n\}$.
-We consider the following three-stage estimation: for each point $x_{0}\in{\mathcal{X}}$ of interest:
+where ${\mathcal{I}}^{(\ell)}$ is the index set of ${\mathcal{D}}^{(\ell)}$, i.e., ${\mathcal{I}}^{(1)}\cup{\mathcal{I}}^{(2)}=\{1,2,\dots,n\}$.
+We consider the following three-stage estimation for each point $x_{0}\in{\mathcal{X}}$ of interest:
 First-stage:
-Estimate $f_{0}$ using any smooth model and denote the estimator by $\widehat{f}_{n}$.
+Estimate $f_{0}$ using any smooth model, and denote this estimator by $\widehat{f}_{n}$.
 Second-stage:
-Estimate the conditional expected residual of the first-stage estimator ${\mathbb{E}}\big{[}Y-f(X)\mid X=x_{0}\big{]}$ or, equivalently, the estimation error $f_{0}(x_{0})-\widehat{f}_{n}(x_{0})$.
-We denote the estimator by
+Estimate the conditional expected residual of the first-stage estimator, ${\mathbb{E}}\big{[}Y-f(X)\mid X=x_{0}\big{]}$, or equivalently, the estimation error $f_{0}(x_{0})-\widehat{f}_{n}(x_{0})$.
+Denote this estimator by $\widehat{b}_{n}(x_{0})$.
 Third-stage:
-Summing the first and second stage estimators and obtain
+Sum the first- and second-stage estimators to obtain
 $\widetilde{f}_{n}(x_{0})\coloneqq\widehat{b}_{n}(x_{0})+\widehat{f}_{n}(x_{0}),$ (1)
-as a debiased estimator of $f_{0}$.
-As explained in Section 3, we can show the asymptotic normality and uniform convergence for the debiased estimator.
-In contrast, many of the modern machine learning methods lack those properties, even though their performances are guaranteed for the population risk.
-In this section, we briefly explain the construction of the first-stage estimator $\widehat{f}_{n}$ and the second-stage estimator $\widehat{b}_{n}$.
+which serves as a debiased estimator of $f_{0}$.
+The term debiased estimator refers to our objective of correcting bias arising from the first-stage estimator.
+In many nonparametric and machine learning methods, bias can be problematic.
+By subtracting the estimated residual, ${\mathbb{E}}\big{[}Y-\widehat{f}_{n}(X)\mid X=x_{0}\big{]}$, we compensate for the difference between $f_{0}$ and $\widehat{f}_{n}$, resulting in an estimator with improved convergence and inferential properties.
+Our estimator (1) is closely related to the influence function in conditional mean estimation (Kennedy, 2023).
+It is also connected to the Neyman orthogonal score (Chernozhukov et al., 2018), which is essentially equivalent to the canonical gradient in one-step bias correction (Schuler & van der Laan, 2024).
+Notably, in conditional mean estimation, constructing the influence function is more intricate due to the conditioning.
+For an influence function that involves parameters represented via the conditional expected value, see Ichimura & Newey (2022).
+Independently and concurrently, Chernozhukov et al.
+(2024) investigates a related topic with different motivations and methods.
+3 The debiased nonparametric regression with local polynomial conditional expected residual estimation
+In this section, we present an example of our debiased estimator.
+We employ local polynomial regression to estimate the conditional expected residual.
+For simplicity, we focus on the case where the covariate is one-dimensional ($d=1$) and ${\mathcal{X}}=[0,1]$.
 3.1 First-stage nonparametric regression
-For the first-stage estimator $\widehat{f}_{n}$, we can use any regression methods if $f_{0}-\widehat{f}_{n}$ is smooth.
-There can be several definitions for the smoothness.
-In this study, for simplicity, we focus on the smoothness in the meaning of the Hölder class (Definition 3.1).
-We note that both $f_{0}$ and $\widehat{f}_{n}$ do not have to be smooth if the difference $f_{0}-\widehat{f}_{n}$ is smooth.
-We emphasize that in our analysis, we do not require specific properties for $\widehat{f}_{n}$ except for the smoothness.
-The first-stage estimator $\widehat{f}_{n}$ can converge to the true function $f_{0}$ with a very slow rate or even be inconsistent.
-Unless the smoothness conditions are satisfied, our asymptotic theoretical results hold, although the empirical finite-sample performance depends on the choice of the first stage estimator.
+For the first-stage estimator $\widehat{f}_{n}$, any regression method can be used, provided that smoothness conditions are met.
+As we show later, our key assumption is that the difference $f_{0}-\widehat{f}_{n}$ is smooth.
+While various notions of smoothness exist, for simplicity, we focus on smoothness in the sense of the Hölder class (Definition 4.1).
+Notably, neither $f_{0}$ nor $\widehat{f}_{n}$ needs to be smooth individually, as long as their difference $f_{0}-\widehat{f}_{n}$ is smooth.
+We emphasize that our analysis does not impose any specific requirements on $\widehat{f}_{n}$ beyond smoothness.
+The first-stage estimator $\widehat{f}_{n}$ may converge to $f_{0}$ at an arbitrarily slow rate or even be inconsistent.
+Nevertheless, as long as the smoothness conditions hold, our asymptotic theoretical results remain valid.
+However, to achieve strong finite-sample performance, a high-quality first-stage estimator is typically preferred.
+In practice, methods such as random forests, neural networks, or other advanced machine learning techniques can be employed, provided they satisfy the smoothness conditions to a reasonable extent.
 3.2 Second-stage conditional expected residual estimation
-For each element $X_{j}$ of covariates $X=(X_{1}\ X_{2}\ \cdots\ X_{d})^{\top}\in{\mathcal{X}}$, let us define the Legendre polynomial series basis $\widetilde{\rho}\colon{\mathbb{R}}\to{\mathbb{R}}$ as
-$\widetilde{\rho}(X_{j})\coloneqq\Big{(}\rho_{0}(X_{j})\ \ \rho_{1}(X_{j})\ \ \cdots\ \ \rho_{\lfloor\ell\rfloor}(X_{j})\Big{)}^{\top},$
-where
-$\widetilde{\rho}_{m}(X_{j})\coloneqq\sum^{m}_{k=0}(-1)^{\ell+m}\sqrt{2m+1}\begin{pmatrix}m\\ k\end{pmatrix}\begin{pmatrix}m+k\\ k\end{pmatrix}X^{k}_{j}.$
-Then, we define $\rho(X_{i})$ to be the corresponding tensor product of all interactions of $\widetilde{\rho}_{m}(X_{1,i}),\dots,\widetilde{\rho}_{m}(X_{d,i})$.
-By using the Legendre polynomial basis, we define the local linear regression as follows:
-$\widehat{\beta}_{n}(x)\coloneqq\begin{pmatrix}\widehat{\beta}_{1,n}(x)&\widehat{\beta}_{2,n}(x)&\cdots&\widehat{\beta}_{\ell,n}(x)\end{pmatrix}^{\top}$
+In the second-stage estimation of the conditional expected residual, we employ local polynomial regression. We define the $\ell$-th order polynomial basis as follows:
+$\rho(u)\coloneqq\begin{pmatrix}1&u&u^{2}/2!&\cdots&u^{\ell}/\ell!\end{pmatrix}^{\top}.$
+In the theoretical analysis, the value of $\ell$ is chosen based on the smoothness of $f_{0}-\widehat{f}_{n}$.
+Using this polynomial basis, we define the local polynomial regression estimator as follows:
+$\widehat{\beta}_{n}(x)\coloneqq\begin{pmatrix}\widehat{\beta}_{0,n}(x)&\widehat{\beta}_{1,n}(x)&\widehat{\beta}_{2,n}(x)&\cdots&\widehat{\beta}_{\ell,n}(x)\end{pmatrix}^{\top}$
 $\coloneqq\operatorname*{arg\,min}_{\beta\in{\mathbb{R}}^{\ell+1}}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})-\beta^{\top}\rho\left(\frac{X_{i}-x}{h_{n}}\right)\right)^{2}K\left(\frac{X_{i}-x}{h_{n}}\right),$
-where $K\colon{\mathcal{X}}\to{\mathbb{R}}$ is a kernel function, defined as
-$K(x)\coloneqq K_{h}(x)\coloneqq\mathbbm{1}\left[\big{\|}x-x_{0}\big{\|}\leq h\right].$
-Then, the conditional expected residual estimator corresponds to $\widehat{\beta}_{1,n}(x)$, the first element of $\widehat{\beta}_{n}(x)$; that is,
-$\widehat{b}_{n}(x)=\widehat{\beta}_{1,n}(x).$
-This second-stage estimator corresponds to the bias-correction term in the construction of semiparametric efficient estimators (Schuler & van der Laan, 2024).
-Such an estimator is referred to as the bias-correction or one-step estimator in the context of semiparametric statistics (van der Vaart, 2002).
-2.3 Discussion
-Our estimator (1) is closely related to the influence function in conditional mean estimation (Kennedy, 2023).
-This type of estimator is also connected to the Neyman orthogonal score (Chernozhukov et al., 2018), which is almost mathematically equivalent to the canonical gradient in the one-step bias corection (Schuler & van der Laan, 2024).
-Note that in conditional mean estimation, the construction of the influence function becomes complicated due to the conditioning.
-For the influence function including the parameter represented via the conditional expected value, see Ichimura & Newey (2022).
-Independently and simultaneously of us, Chernozhukov et al.
-(2024) investigates a related topic, with different interest and method.
+where $K\colon{\mathcal{X}}\to{\mathbb{R}}$ is a kernel function defined by
+$K(u)\coloneqq K_{h}(u)\coloneqq\mathbbm{1}\left[\big{|}u\big{|}\leq h\right].$
+The conditional expected residual estimator is then given by the first component of $\widehat{\beta}_{n}(x)$:
+$\widehat{b}_{n}(x)=\widehat{\beta}_{0,n}(x).$
 4 Convergence analysis
-This section presents convergence analysis of $\widetilde{f}_{n}$. First, we define the Hölder class as follows:
+This section presents a convergence analysis of $\widetilde{f}_{n}$ defined in Section 3. We begin by defining the Hölder class.
 Definition 4.1 (Hölder class).
-Let ${\mathcal{T}}$ be an interval in ${\mathbb{R}}^{d}$ and let $s$ and $L$ be two positive numbers.
-Let $D^{\alpha}=\frac{\partial^{\alpha}}{\partial x_{1}^{\alpha_{1}}\dots\partial x_{d}^{\alpha_{d}}}$ be the partial derivative operator.
-The Hölder class $\Sigma(s,L)$ on ${\mathcal{T}}$ is defined as the set of $\ell=\lfloor s\rfloor$ times differentiable function $f\colon{\mathcal{T}}\to{\mathbb{R}}$ satisfying the following conditions:
-  • Its derivatives up to order $\lfloor\ell\rfloor$ are bounded as
-    $\left|D^{\alpha}f(x)\right|\leq C<\infty\qquad\forall x\in\mathcal{X},$
-    for all $\alpha=(\alpha_{1},\dots,\alpha_{d})$ with $\sum_{j}\alpha_{j}\leq\ell$.
-  • Its $\ell$-order derivatives satisfying the Lipschitz condition: there exists a constant $0<C\infty$ such that
-    $\left|D^{\beta}f(x)-D^{\beta}f(x^{\prime})\right|\leq C\|x-x^{\prime}\|^{s-\ell}\qquad\forall x,x^{\prime}\in\mathcal{X}$
-    for all $\beta=(\beta_{1},\dots,\beta_{d})$ with $\sum_{j}\beta_{j}=\ell$.
-Define
-^$\widehat{{\mathcal{B}}}_{n}(x_{0})$ $\coloneqq\frac{1}{mh^{d}_{n}}\sum_{i\in{\mathcal{I}}^{(2)}}\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{X_{i}-x_{0}}{h_{n}}\right),$
-$\widehat{w}_{h}(x,x_{0})$ $\coloneqq\frac{1}{h^{d}_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(\frac{x-x_{0}}{h_{n}}\right)K\left(\frac{x-x_{0}}{h_{n}}\right).$
-Then, the second-stage estimator is given as
-$\widehat{b}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0}).$
+Given an interval ${\mathcal{T}}\subset{\mathbb{R}}$, and positive constants $\beta$ and $L$, the Hölder class $\Sigma(\beta,L)$ on ${\mathcal{T}}$ is the set of $\ell=\lfloor\beta\rfloor$ times differentiable functions $f\colon{\mathcal{T}}\to{\mathbb{R}}$ whose $\ell$-times derivative $f^{(\ell)}$ satisfies
+$\Big{|}f^{(\ell)}(x)-f^{(\ell)}(z)\Big{|}\leq L\big{|}x-z\big{|}^{\beta-\ell}\quad\forall x,z\in{\mathcal{T}}.$
+4.1 Closed-form solution
+The debiased estimator with local polynomial regression admits a closed-form solution. We define
+^$\widehat{{\mathcal{B}}}_{n}(x_{0})$ $\coloneqq\frac{1}{mh_{n}}\sum_{i\in{\mathcal{I}}^{(2)}}\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)\rho\left(\frac{X_{i}-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{X_{i}-x_{0}}{h_{n}}\right),$
+$\widehat{w}_{h}(x,x_{0})$ $\coloneqq\frac{1}{h_{n}}\rho(0)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(\frac{x-x_{0}}{h_{n}}\right)K\left(\frac{x-x_{0}}{h_{n}}\right).$
+Then, the second-stage estimator is given by
+$\widehat{b}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})\right)\widehat{w}_{h}(X_{i},x_{0}).$
+Thus, the debiased estimator is expressed as
+$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})\right)\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$
 4.2 Bias and variance decomposition
-First, we investigate the bias and variance of $\widetilde{f}_{n}$. That is, we aim to analyze the bias and variance of $\widetilde{b}_{n}(x_{0})$, denoted as
+We begin by analyzing the bias and variance of $\widetilde{f}_{n}$, defined as
 $\mathrm{Bias}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\right]-f_{0}(x_{0}),$
-$\mathrm{Variance}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\right]\Big{)}^{2}\right].$
-Let us introduce the population values of ^$\widehat{{\mathcal{B}}}_{n,x}$ and $\widehat{s}_{n,x}$ as
+$\mathrm{Variance}(x_{0})$ $\coloneqq{\mathbb{E}}\left[\left(\widetilde{f}_{n}(x_{0})-{\mathbb{E}}[\widetilde{f}_{n}(x_{0})]\right)^{2}\right].$
+The bias term quantifies how closely the estimator $\widetilde{f}_{n}$ centers around the true regression function $f_{0}$ at $x_{0}$, while the variance term measures the fluctuation of $\widetilde{f}_{n}$ around its own mean.
+Balancing bias and variance is a fundamental concern in statistical estimation.
+Our approach, which employs local polynomial smoothing in the second stage, aims to systematically reduce the bias introduced by the first-stage estimator while controlling variance through appropriate bandwidth selection.
+To facilitate our analysis, we introduce the population version of ^$\widehat{{\mathcal{B}}}_{n}(x_{0})$:
 ~$\widetilde{{\mathcal{B}}}_{n}(x_{0})\coloneqq\int\rho\left(\frac{x-x_{0}}{h_{n}}\right)\rho\left(\frac{x-x_{0}}{h_{n}}\right)^{\top}K\left(\frac{x-x_{0}}{h_{n}}\right){\mathrm{d}}F(x).$
-For this value, the covariate distribution, and the smoothness of $f_{0}$ and $\widehat{f}_{n}$, we make the following assumptions.
-Assumption 3.2.
-Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For a point $x_{0}\in{\mathcal{X}}$ of interest, the following holds:
-  • $f_{0}-\widehat{f}$ belongs to the Hölder class $\Sigma(s,L)$.
-  • The eigenvalues of ^$\widehat{{\mathcal{B}}}_{n}(x)$ is bounded above and below away from zero.
-  • There exists a constant $C>0$ such that $\int\mathbbm{1}\big{[}\|x-x_{0}\|\leq h\big{]}{\mathrm{d}}P(x)\leq Ch^{d}$,
-where $P(x)$ denotes a distribution of $x$.
+We then establish bounds for the bias and variance terms as follows:
 Theorem 4.2 (Bias and variance decomposition).
-Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$.
-For the values, suppose that Assumptions 3.2 holds.
-Let $h_{n}$ be the bandwidth of the local polynomial estimator.
-Then, for any $\varepsilon>0$ and for all $x_{0}\in[0,1]$, there exits $n_{0}>0$ such that for all $n\geq n_{0}$, with probability $1-\varepsilon$, it holds that
-$\left|\mathrm{Bias}(x_{0})\right|\leq C_{1}h^{s}_{n},$
-$\mathrm{Variance}(x_{0})\leq\frac{C_{2}}{nh^{d}_{n}}+C_{3}h^{2s}_{n},$
-where $C_{1},C_{2},C_{3}>0$ are constants independent of $f_{0}$ and $n$.
-The smoothness condition on $f_{0}-\widehat{f}$ can be satisfied in several ways. We discuss this point in Section 4.
-3.2 Pointwise convergence of the mean squared errors
-Theorem 3.3 about the convergences of bias and variance terms directly indicate the convergence of the mean squared errors. This is because we have
-${\mathbb{E}}\left[\left(\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=\mathrm{Bias}(x_{0})^{2}+\mathrm{Variance}(x_{0})$
-To minimize the mean squared error, we set the bandwidth as
-$h_{n}=\alpha n^{-\frac{1}{2s+d}},$
-where $\alpha$ is a constant independent of $f_{0}$ and $n$, and recall that $s$ is the parameter of smoothness and $d$ is the dimension of $X$.
-Then, we state the pointwise convergence of the mean squared errors as follows:
-Theorem 3.4 (Pointwise mean squared error convergence).
-Suppose that the conditions in Theorem 3.3 holds. Let $h_{n}=\alpha n^{-\frac{1}{2s+d}}$. Then, for all $x_{0}\in[0,1]$, it holds that
-$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}\sup_{x_{0}\in{\mathcal{X}}}{\mathbb{E}}_{f}\left[\psi^{-2}_{n}\Big{|}\widetilde{f}_{n}(x_{0})-f(x_{0})\Big{|}^{2}\right]\leq C<\infty,$
-where $\psi_{n}\coloneqq n^{-\frac{s}{2s+d}}$ is the rate of convergence and $C$ is a constant depending only on $s$, $L$, $\lambda_{0}$, $a_{0}$, $\sigma^{2}_{\max}$, $K_{\max}$, and $s$.
-Theorem 3.4 can be extend to the following mean squared error convergence over the distribution of $X$.
-Corollary 3.5 (Mean squared errors over the distribution of $X$).
-Suppose that the conditions in Theorem 3.4 holds. Then, it holds that
-$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}{\mathbb{E}}_{f}\left[\psi^{-2}_{n}\Big{\|}\widetilde{f}_{n}-f\Big{\|}^{2}_{2}\right]\leq C<\infty,$
-where $\psi_{n}\coloneqq n^{-\frac{s}{2s+d}}$ is the rate of convergence and $C$ is a constant depending only on $s$, $L$, $\lambda_{0}$, $a_{0}$, $\sigma^{2}_{\max}$, $K_{\max}$, and $s$.
+Let $s,L,C,C_{1},C_{2},C_{3}>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For every $x\in{\mathcal{X}}$, the following hold:
+  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
+  • There exists a constant $C>0$ such that
+    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
+    where $P(x)$ denotes the distribution of $x$.
+If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then for any $\varepsilon>0$ and for all $x_{0}\in{\mathcal{X}}$, there exists $n_{0}>0$ such that for all $n\geq n_{0}$, with probability at least $1-\varepsilon$, it holds that
+$\bigl{|}\mathrm{Bias}(x_{0})\bigr{|}$ $\leq C_{1}h^{s}_{n},$
+$\mathrm{Variance}(x_{0})$ $\leq\frac{C_{2}}{nh_{n}}+C_{3}h^{2s}_{n}.$
+The smoothness condition on $f_{0}-\widehat{f}_{n}$ can be satisfied in various ways; see Section 5 for further discussion.
+The above result implies that as long as the smoothness of $f_{0}-\widehat{f}_{n}$ is well-controlled and the bandwidth $h_{n}$ is chosen appropriately to shrink with $n$, both bias and variance remain bounded and ultimately vanish as $n$ grows.
+4.3 Pointwise convergence of the MSE
+From Theorem 4.2, the bounds on bias and variance imply the convergence of the MSE, noting that
+${\mathbb{E}}\left[\left(\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=\mathrm{Bias}(x_{0})^{2}+\mathrm{Variance}(x_{0}).$
+To minimize the MSE, we set
+$h_{n}=\alpha n^{-\frac{1}{2s+1}},$
+where $\alpha$ is a constant independent of $f_{0}$ and $n$, $s$ is the smoothness parameter, and $d$ is the dimension of $X$. We then establish the pointwise MSE convergence result.
+Theorem 4.3 (Pointwise MSE convergence).
+Let $s,L,C,\alpha>0$ be constants independent of $f_{0}$ and $n$. Let
+$h_{n}=\alpha n^{-\frac{1}{2s+1}}$
+be the bandwidth of the local polynomial estimator. For every $x_{0}\in{\mathcal{X}}$, the following hold:
+  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
+  • There exists a constant $C>0$ such that
+    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
+    where $P(x)$ denotes the distribution of $x$.
+If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then for all $x_{0}\in{\mathcal{X}}$, it holds that
+${\mathbb{E}}\left[\left(\widetilde{f}_{n}(x_{0})-f_{0}(x_{0})\right)^{2}\right]=O\left(n^{-\frac{s}{2s+1}}\right),$
+as $n\to\infty$.
+Theorem 4.3 can be extended to establish the convergence of the MSE over the distribution of $X$.
+Corollary 4.4 (MSE over the distribution of $X$).
+Suppose the conditions in Theorem 4.3 hold. Then, we have
+${\mathbb{E}}\left[\left(\widetilde{f}_{n}(X)-f_{0}(X)\right)^{2}\right]=O\left(n^{-\frac{2s}{2s+1}}\right),$
+where the expectation is taken over both $\widetilde{f}_{n}$ and $X$, as $n\to\infty$.
 4.4 Pointwise asymptotic normality
@@
 Theorem 4.5 (Asymptotic normality).
-Suppose one of the upper bounds in Theorem 3.4 holds. Then, it tholds that
-$\sqrt{nh^{d}_{n}}\Big{(}\widetilde{f}(x_{0})-f_{0}(x_{0})\Big{)}\xrightarrow{{\mathrm{d}}}{\mathcal{N}}(0,V(x_{0})),$
-where $V(x_{0})\coloneqq nh^{d}_{n}\mathrm{Variance}(x_{0})\leq C_{2}+o(h_{n})$.
-The asymptotic normality result facilitates statistical inference.
-However, constructing a confidence interval is challenging due to the unknown parameter $s$ involved in $h_{n}$.
-To address this issue, the concept of honest confidence intervals has been developed.
-For example, the bootstrap method can be employed to obtain valid confidence intervals, even when certain parameters remain unspecified.
+Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$.
+Let $h_{n}$ be the bandwidth of the local polynomial estimator such that $h_{n}\to 0$, $nh_{n}\to\infty$, and $nh^{2s+1}_{n}\to 0$ hold as $n\to\infty$.
+For every $x\in{\mathcal{X}}$, the following hold:
+  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
+  • There exists a constant $C>0$ such that
+    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
+    where $P(x)$ is the distribution of $x$.
+If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then for every $x_{0}\in{\mathcal{X}}$, it holds that
+$\sqrt{nh_{n}}\left(\widetilde{f}(x_{0})-f_{0}(x_{0})\right)\xrightarrow{{\mathrm{d}}}{\mathcal{N}}\big{(}0,V(x_{0})\big{)},$
+as $n\to\infty$, where $V(x_{0})\geq 0$ is a constant depending on $x_{0}$ but independent of $n$.
+Here, note that the following holds from the bias and variance decomposition:
+$V(x_{0})\coloneqq nh_{n}\mathrm{Variance}(x_{0})\leq C_{2}+o(h_{n}).$
+The asymptotic normality result facilitates statistical inference, such as constructing confidence intervals or hypothesis tests for $f_{0}(x_{0})$.
+In practice, knowing that $\sqrt{nh_{n}}(\widetilde{f}(x_{0})-f_{0}(x_{0}))$ converges in distribution to a normal allows for classical inferential techniques, provided that the variance $V(x_{0})$ can be consistently estimated.
+However, constructing a confidence interval can be challenging, because $s$ (the smoothness parameter) is often unknown and affects the choice of $h_{n}$.
+In real applications, data-driven methods, including cross-validation, can be employed to select $h_{n}$, and bootstrap or other resampling approaches can help in variance estimation.
 4.5 Uniform convergence
-Lastly, we establish the uniform (or sup-norm) convergence of $\widetilde{f}(x_{0})$.
-Uniform convergence bounds the supremum of the estimation error uniformly over $X\in{\mathcal{X}}$, without relying on the distribution of $X$, whereas mean squared convergence depends on the distribution of $X$.
-This theoretical guarantee is particularly useful for addressing distribution shift problems, such as covariate shift (Shimodaira, 2000).
-In covariate shift scenarios, the distribution of $X$ differs between the training and test data.
-While mean squared convergence may be affected by such a shift, uniform convergence remains unaffected, ensuring robust performance for both training and test data.
-Consequently, uniform convergence is critical in discussions involving distribution shifts.
+Finally, we establish the uniform (sup-norm) convergence of $\widetilde{f}_{n}(x_{0})$.
+Unlike mean squared convergence, which depends on the distribution of $X$, uniform convergence bounds the supremum of the estimation error over all $x\in{\mathcal{X}}$.
+This property is particularly valuable for addressing the covariate shift problem (Shimodaira, 2000; Schmidt-Hieber & Zamolodtchikov, 2024), as uniform convergence remains unaffected by changes in the distribution of $X$ between training and test datasets.
 Theorem 4.6 (Uniform convergence).
-Suppose that $f_{0}$ belongs to a Hölder class $\Sigma(s,L)$ on $\mathcal{X}$, where $s>0$ and $L>0$. Let $\widehat{f}_{n}$ be the $\ell$) estimator of order $\ell=\lfloor s\rfloor$ with bandwidth
-$h_{n}\coloneqq\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+d}}$
-for some $\alpha>0$. Suppose the following:
-  • Assumption 3.2 holds.
-  • $\varepsilon_{i}$ is a sub-Gaussian random variable satisfying
-    ${\mathbb{E}}\big{[}\exp\big{(}\lambda\varepsilon_{i}\big{)}\big{]}\leq\exp\big{(}K^{2}\lambda^{2}\big{)},\quad\forall\lambda\in{\mathbb{R}},$
+Let $s,L_{K},L,C,\alpha>0$ be constants independent of $f_{0}$ and $n$. Let
+$h_{n}\coloneqq\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+1}}$
+be the bandwidth of the local polynomial estimator. For every $x_{0}\in{\mathcal{X}}$, the following hold:
+  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
+  • There exists a constant $C>0$ such that
+    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
+    where $P(x)$ denotes the distribution of $x$.
+  • The noise $\varepsilon_{i}$ is sub-Gaussian, satisfying
+    ${\mathbb{E}}\left[\exp(\lambda\varepsilon_{i})\right]\leq\exp\left(K^{2}\lambda^{2}\right)\quad\forall\lambda\in{\mathbb{R}},$
     for some constant $K>0$.
-  • $K$ is a Lipschitz kernel: $K\in\Sigma(1,L_{K})$ on $\cal{X}$ with $0<L_{K}<\infty$.
-Then, there exists a constant $C<\infty$ such that
-$\limsup_{n\to\infty}\sup_{f\in\Sigma(s,L)}{\mathbb{E}}_{f}\big{[}\psi^{-2}_{n}\big{\|}\widetilde{f}_{n}-f\big{\|}^{2}_{\infty}\big{]}\leq C,$
-where
-$\psi_{n}\coloneqq\left(\frac{\log(n)}{n}\right)^{\frac{s}{2s+d}}.$
-Uniform convergence provides robustness against distributional shifts but has not been proven for many nonparametric regression estimators produced by machine learning methods.
-This limitation arises partly from their data-adaptive nature and the reliance on empirical process techniques within the empirical risk minimization framework (van de Geer, 2009).
-Using empirical processes and model complexity, estimation error can be evaluated in a general manner, avoiding significant dependence on specific algorithms.
-However, such approaches often focus on population risk, such as mean squared error, without addressing how these results translate into uniform convergence guarantees.
-For example, Schmidt-Hieber & Zamolodtchikov (2024) tackle this issue by proposing an estimator based on neural networks, demonstrating restricted uniform convergence in the case of one-dimensional covariates.
-In contrast, our proposed method is model-free and accommodates multiple-dimensional covariates, offering broader applicability and robustness.
+  • The kernel function $K$ is Lipschitz: $K\in\Sigma(1,L_{K})$ on ${\mathcal{X}}$.
+If $f_{0}-\widehat{f}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, then the following holds:
+${\mathbb{E}}\left[\left\|\widetilde{f}_{n}-f_{0}\right\|^{2}_{\infty}\right]=O\left(\frac{\log(n)}{n}\right)^{\frac{2s}{2s+1}}.$
+Although uniform convergence ensures robust performance under distribution shifts, it has not been established for many nonparametric regression estimators derived from modern machine learning methods.
+One reason for this gap is their data-adaptive nature and the reliance on empirical process techniques, which primarily focus on population risk measures whose expectation is taken over the covariate distribution rather than on sup-norm bounds.
+For example, Schmidt-Hieber & Zamolodtchikov (2024) analyzes a neural network-based estimator in a one-dimensional covariate setting, demonstrating restricted uniform optimality.
+In contrast, our approach is model-free and accommodates multi-dimensional covariates, offering broader applicability and greater robustness.
 4.6 Double robustness
-Our proposed debiased estimator possesses the property of double robustness; that is, if either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, the resulting debiased estimator $\widetilde{f}_{n}$ is also consistent.
-Recall that the debiased estimator is defined as
-$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$
-Suppose that $\widehat{f}_{n}(x_{0})$ converges in probability to $f^{\dagger}(x_{0})$ and $\widehat{b}_{n}(x_{0})$ converges in probability to $b^{\dagger}(x_{0})$.
-We examine the behavior of $\widetilde{f}_{n}(x_{0})$ under two cases: $f^{\dagger}(x_{0})=f_{0}(x_{0})$ and $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$.
-First, consider the case where $f^{\dagger}(x_{0})=f_{0}(x_{0})$ and $b^{\dagger}(x_{0})\neq f_{0}(x_{0})-f^{\dagger}(x_{0})$. Then, it follows that
+Our proposed debiased estimator also possesses the property of double robustness: if either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, then $\widetilde{f}_{n}$ remains consistent.
+Recall that
+$\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\left(Y_{i}-\widehat{f}_{n}(X_{i})\right)\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0}).$
+Suppose $\widehat{f}_{n}(x_{0})$ converges in probability to $f^{\dagger}(x_{0})$ and $\widehat{b}_{n}(x_{0})$ converges in probability to $b^{\dagger}(x_{0})$. We analyze two cases:
+  • If $f^{\dagger}(x_{0})=f_{0}(x_{0})$ but $b^{\dagger}(x_{0})\neq f_{0}(x_{0})-f^{\dagger}(x_{0})$, then
     $\widetilde{f}_{n}(x_{0})=\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-f_{0}(X_{i})\Big{)}w^{\dagger}(X_{i},x_{0})+f_{0}(x_{0})+o_{p}(1),$
-where $w^{\dagger}(X_{i},x_{0})$ is a function depending only on $X_{i}$ and $x_{0}$.
-Since the first term, $\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}Y_{i}-f_{0}(X_{i})\Big{)}w^{\dagger}(X_{i},x_{0})$, converges in probability to zero, $\widetilde{f}_{n}(x_{0})$ converges in probability to $f_{0}(x_{0})$.
-Next, consider the case where $f^{\dagger}(x_{0})\neq f_{0}(x_{0})$ and $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$. Then, it holds that
-$\widetilde{f}_{n}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})+f^{\dagger}(x_{0})+o_{p}(1)=f_{0}(x_{0})+o_{p}(1).$
-Thus, in either case, $\widetilde{f}_{n}(x_{0})$ converges in probability to $f_{0}(x_{0})$. We summarize this property in the following theorem.
+    where $w^{\dagger}(X_{i},x_{0})$ depends only on $X_{i}$ and $x_{0}$.
+Since the first term converges in probability to zero, it follows that $\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})$.
+  • If $f^{\dagger}(x_{0})\neq f_{0}(x_{0})$ but $b^{\dagger}(x_{0})=f_{0}(x_{0})-f^{\dagger}(x_{0})$, then
+    $\widetilde{f}_{n}(x_{0})=\left(f_{0}(x_{0})-f^{\dagger}(x_{0})\right)+f^{\dagger}(x_{0})+o_{p}(1)=f_{0}(x_{0})+o_{p}(1).$
+In both cases, $\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})$. We summarize this result as follows.
 Theorem 4.7 (Double robustness).
-If either $\widehat{b}_{n}$ or $\widehat{f}_{n}$ is consistent, then it holds that
-$\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})\qquad\mathrm{as}\ \ n\to\infty.$
+For any $x_{0}\in{\mathcal{X}}$, if either $\widehat{b}_{n}(x_{0})$ or $\widehat{f}_{n}(x_{0})$ is consistent, then
+$\widetilde{f}_{n}(x_{0})\xrightarrow{{\mathrm{p}}}f_{0}(x_{0})\qquad\text{as }n\to\infty.$
+Double robustness thus provides an additional safeguard against model misspecification or suboptimal performance of one of the estimators.
+In practice, this property is particularly valuable when the first-stage model $\widehat{f}_{n}$ fails to accurately capture the true function, but the second-stage residual model $\widehat{b}_{n}$ remains reliable—or vice versa.
 5 Smoothness is all you need
-As demonstrated above, our method requires only the smoothness of $\widehat{f}-f_{0}$.
-This assumption is satisfied when $f_{0}$ and $\widehat{f}$ are individually smooth, but it also holds in cases where $f_{0}$ and $\widehat{f}$ are not smooth, provided that their difference is smooth.
-Neural networks belong to this class under the analysis of (Schmidt-Hieber, 2020).
-Remarkably, our results require only the smoothness of the first-stage estimator (the difference between $\widehat{f}_{n}$ and $f_{0}$).
+As demonstrated above, our method relies solely on the smoothness of $\widehat{f}-f_{0}$.
+This assumption holds when both $f_{0}$ and $\widehat{f}$ are individually smooth, but it also remains valid in cases where neither $f_{0}$ nor $\widehat{f}$ is smooth, as long as their difference is smooth.
+Neural networks satisfy this condition under the analysis in Schmidt-Hieber (2020).
+Notably, our results require only the smoothness of the first-stage estimator (i.e., the difference $\widehat{f}_{n}-f_{0}$).
 This assumption significantly broadens the applicability of our method.
-Even when $f_{0}$ and $\widehat{f}_{n}$ are not smooth, our method can still guarantee mean squared convergence, asymptotic normality, and uniform convergence.
-For instance, Imaizumi & Fukumizu (2019) discusses the convergence of nonparametric estimators using neural networks when $f_{0}$ belongs to a piecewise Hölder class.
-Under such a case, the liner estimators, including locally polynomial regression, are suboptimal.
-However, even under the case, the difference $\widehat{f}-f_{0}$ can be smooth, and our smoothness condition still holds.
+Even when $f_{0}$ and $\widehat{f}_{n}$ are not individually smooth, our approach can still ensure MSE convergence, asymptotic normality, and uniform convergence, provided that the difference is smooth.
+From a different perspective, this flexibility allows one to use state-of-the-art or specialized methods in the first stage without compromising the theoretical guarantees of the final estimator, as long as the first-stage error exhibits a level of smoothness akin to the Hölder condition.
+This feature is particularly useful when the true regression function $f_{0}$ has a complex or partially known structure, yet the difference $f_{0}-\widehat{f}_{n}$ remains more regular due to design choices or inductive biases in $\widehat{f}_{n}$.
 6 Conclusion
 This study developed a debiased estimator for nonparametric regression.
-For any smooth regression estimator, our debiased estimator gives the following desirable property: mean-squared convergence, asymptotic normality, uniform convergence, and double robustness.
-The asymptotic normality is useful for statistical inference, while the uniform convergence makes the nonparametric regression estimator distributionally robust.
-Additionally, either the first-stage or the second-stage nonparametric estimator is consistent, the resulting regression estimator is consistent; that is, doubly robust.
-Surprisingly, our estimator only requires the smoothness of the first-stage regression estimator.
-More specifically, if either (i) both the first-stage estimator $\widehat{f}$ and the true regression function $f_{0}$ are smooth or (ii) the difference $\widehat{f}-f_{0}$ is smooth, we can show the above desirable properties for the debiased estimator.
-6 Proofs
+For any smooth regression estimator, our debiased approach ensures pointwise and uniform MSE convergence, asymptotic normality, and double robustness.
+The asymptotic normality facilitates statistical inference, while uniform convergence provides distributional robustness for nonparametric regression.
+Additionally, if either the first-stage or the second-stage nonparametric estimator is consistent, then the resulting regression estimator remains consistent, exhibiting the doubly robust property.
+Notably, our estimator requires only the smoothness of the first-stage regression estimator.
+Specifically, if either (a) both the first-stage estimator $\widehat{f}$ and the true regression function $f_{0}$ are smooth, or (b) the difference $\widehat{f}-f_{0}$ is smooth, then the desirable properties of the debiased estimator hold.
+Thus, our method imposes minimal structural constraints while guaranteeing strong theoretical results, making it an attractive and flexible approach for a wide range of nonparametric regression problems.
+Appendix
+Appendix A Proofs
 This section provides the proofs in Section 4.
-6.1 Proof of Theorem 3.6
+A.1 Proof of Theorem 4.5
 Define
@@
 $\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\coloneqq{\mathbb{E}}\left[\Big{(}\widetilde{f}_{n}(x_{0})-{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{)}^{2}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right],$
-$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$ $\coloneqq{\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}.$
+$\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
+$\ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \ \coloneqq{\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}.$
 Here, the followings hold:
 $\mathrm{Bias}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Bias}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right],$
-$\mathrm{Variance}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right]+\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right).$
-We aim to show that
+$\mathrm{Variance}\left(x_{0}\right)={\mathbb{E}}\left[\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\right]+\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right).$
+We aim to show that the followings hold with probability one:
 $\mathrm{Bias}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)\leq\frac{L}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(X_{i},x_{0})\Big{|},$ (2)
-$\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)=\frac{C_{2}}{nh^{d}_{n}}+o(1),$ (3)
-$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)=2C^{2}_{1}h^{2s}.$ (4)
+$\mathrm{Variance}\left(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right)=\frac{C_{2}}{nh_{n}}+o(1),$ (3)
+$\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)=2C^{2}_{1}h^{2s}.$ (4)
 If (2) holds, then we have
 ${\mathbb{E}}\left[\frac{L}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(X,x_{0})\Big{|}\right]$ $=\int L\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(x,x_{0})\Big{|}{\mathrm{d}}P(x){\mathrm{d}}x$
-$=\int L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u,$
+$=\int L\frac{h^{s}_{n}}{\ell!}h_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u,$
 where $\frac{x-x_{0}}{h}=u$. We have
-$\int L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u$
-$=L\frac{h^{s}_{n}}{\ell!}h^{d}_{n}\int\left|\frac{1}{h^{d}_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(u\right)\mathbbm{1}\left[\|u\|\leq 1\right]\right|{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u.$
-6.1.1 Preliminaries
+$\int L\frac{h^{s}_{n}}{\ell!}h_{n}\Big{|}\widehat{w}_{h}(x_{0}+h_{n}u,x_{0})\Big{|}{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u$
+$=L\frac{h^{s}_{n}}{\ell!}h_{n}\int\left|\frac{1}{h_{n}}\rho\left(0\right)^{\top}\widehat{{\mathcal{B}}}_{n}(x_{0})^{-1}\rho\left(u\right)\mathbbm{1}\left[\|u\|\leq 1\right]\right|{\mathrm{d}}P(x_{0}+h_{n}u){\mathrm{d}}u.$
+A.1.1 Preliminaries
 As preliminary, we first state the following lemma about the properties of the local polynomial estimator, as well as Proposition 1.12 in Tsybakov (2008) and Appendix B.3 in Kennedy et al. (2024).
-Lemma 6.1 (From Proposition 1.12 in Tsybakov (2008)).
+Lemma A.1 (From Proposition 1.12 in Tsybakov (2008)).
 Let $x_{0}$ be a real number such that ${\mathcal{B}}_{n}(x_{0})>0$ and let $Q$ be a polynomial whose degree is less than or equal to $\ell$.
@@
 for any sample $(X_{1},\dots,X_{n})$.
-Lemma 6.2.
-Under Assumption 3.2, for any $\varepsilon$, there exists $n_{0}$ such that for all $n\geq n_{0}$, $h\geq\frac{1}{2n}$, and $x_{0}\in\mathcal{X}$, the weights $\widehat{w}_{n}(x,x_{0})$ of the $\ell$) estimator satisfy the followings:
+Lemma A.2.
+Let $s,L,C>0$ be constants independent of $f_{0}$ and $n$. Let $h_{n}$ be the bandwidth of the local polynomial estimator. For every $x_{0}\in{\mathcal{X}}$, the following hold:
+  • The eigenvalues of ~$\widetilde{{\mathcal{B}}}_{n}(x)$ are uniformly bounded above and below away from zero.
+  • There exists a constant $C>0$ such that
+    $\int\mathbbm{1}\big{[}|x-x_{0}|\leq h_{n}\big{]}{\mathrm{d}}P(x)\leq Ch_{n},$
+    where $P(x)$ is the distribution of $x$.
+For any $\varepsilon$, there exists $n_{0}$ such that for all $n\geq n_{0}$, $h\geq\frac{1}{2n}$, and $x_{0}\in\mathcal{X}$, the weights $\widehat{w}_{n}(x,x_{0})$ of the $\ell$) estimator satisfy the followings:
   • $\sup_{x,x_{0}}\Big{|}\widehat{w}_{h}(x,x_{0})\Big{|}\leq\frac{C_{*}}{nh}$ with probability $1-\varepsilon$;
@@
 where the constant $C_{*}$ depends only on $\lambda_{0}$, $a_{0}$, and $K_{\max}$.
-6.1.2 Bounding the bias (proof of (2))
+A.1.2 Bounding the bias (proof of (2))
 Proof.
@@
 $={\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}f_{0}(X_{0})-\widehat{f}_{n}(X_{i})\Big{)}\widehat{w}_{h}(X_{i},x_{0})+\widehat{f}_{n}(x_{0})-f_{0}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right].$
-Let $\widehat{g}_{n}\coloneqq\widehat{f}_{n}-f_{0}$. Since $\widehat{g}_{n}$ belongs to the Hölder class $\Sigma(s,L)$, we have
+Let $\widehat{g}_{n}\coloneqq\widehat{f}_{n}-f_{0}$. Since $\widehat{g}_{n}$ belongs to the Hölder class $\Sigma(s,L)$ almost surely as $n\to\infty$, we have
 ${\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\Big{(}\widehat{g}_{n}(X_{i})-\widehat{g}_{n}(x_{0})\Big{)}\widehat{w}_{h}(X_{i},x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
 $={\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{\widehat{g}^{(\ell)}(x_{0}+\tau_{i}(X_{i}-x_{0}))-\widehat{g}^{(\ell)}(x_{0})}{\ell!}\big{(}X_{i}-x_{0}\big{)}^{\ell}\widehat{w}_{h}(X_{i},x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right].$
-where $0\leq\tau_{i}\leq 1$. Here, we used the Taylor expansion of $\widehat{g}$, and Lemma 6.1.
-With probability one, we have
+as $n\to\infty$, where $0\leq\tau_{i}\leq 1$. Here, we used the Taylor expansion of $\widehat{g}$, and Lemma A.1.
+Therefore, the followng holds almost surely as $n\to\infty$:
 ${\mathbb{E}}\left[\frac{1}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{\widehat{g}^{(\ell)}(x_{0}+\tau_{i}(X_{i}-x_{0}))-\widehat{g}^{(\ell)}(x_{0})}{\ell!}\big{(}X_{i}-x_{0}\big{)}^{\ell}\widehat{w}_{h}(X_{i},x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
@@
 $\leq\frac{L}{m}\sum_{i\in{\mathcal{I}}^{(2)}}\frac{h^{s}_{n}}{\ell!}\Big{|}\widehat{w}_{h}(X_{i},x_{0})\Big{|}.$
-6.1.3 Bounding the variance: part I (proof of (3))
+A.1.3 Bounding the variance: part I (proof of (3))
 Proof.
-We have
+As $n\to\infty$, the following holds almost surely:
 $\mathrm{Variance}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})$
@@
 $=\sum_{i\in{\mathcal{I}}^{(2)}}\left(\widehat{w}_{h}(X_{i},x_{0})\right)^{2}{\mathbb{E}}\big{[}\xi^{2}_{i}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\big{]}.$
-From Assumption 3.2, we have
+From the conditions we assumed, as $n\to\infty$, the following holds:
 ${\mathbb{E}}\left[\mathrm{Variance}(x_{0}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n})\right]$
@@
 $\leq\sigma^{2}_{\max}\frac{1}{m}{\mathbb{E}}\left[\sup_{x,x_{0}}\Big{|}\widehat{w}_{h}(x,x_{0})\Big{|}\right]{\mathbb{E}}\left[\Big{|}\widehat{w}_{h}(X,x_{0})\Big{|}\right]$
-$\leq\frac{\sigma^{2}_{\max}C^{2}_{*}}{mh^{d}_{n}}+o(1)$
-$=\frac{C_{2}}{nh^{d}_{n}}+o(1).$
-6.1.4 Bounding the variance: part II (proof of (4))
+$\leq\frac{\sigma^{2}_{\max}C^{2}_{*}}{mh_{n}}+o(1)$
+$=\frac{C_{2}}{nh_{n}}+o(1).$
+A.1.4 Bounding the variance: part II (proof of (4))
 Proof.
-We have
-$\mathrm{Variance}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
+As $n\to\infty$, the following holds t
+$\mathrm{Var}\left({\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right)$
 $={\mathbb{E}}\Big{[}\Big{(}{\mathbb{E}}\Big{[}\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\Big{]}-f_{0}(x_{0})+f_{0}(x_{0})-{\mathbb{E}}\left[{\mathbb{E}}\left[\widetilde{f}_{n}(x_{0})\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\right]\Big{)}^{2}\Big{]}$
@@
 $\leq 2C^{2}_{1}h^{2s}_{n}.$
-6.2 Proof of Theorem 3.7
+A.2 Proof of Theorem 4.6
 Proof.
-We decompose the sup-norm as follows:
+As $n\to\infty$, the following holds almost surely:
 ${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-f_{0}\Big{\|}^{2}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]$
@@
 Finally, we have
-${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh^{d}_{n}},$
+${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-{\mathbb{E}}\left[\widetilde{f}_{n}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\Big{\|}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh_{n}},$
 where $C_{3}>0$ is a constant independent of $f$ and $n$.
-We have
+As $n\to\infty$, the following holds almost surely:
 ${\mathbb{E}}\left[\Big{\|}\widetilde{f}_{n}-f_{0}\Big{\|}^{2}_{\infty}\mid\{X_{i}\}_{i\in{\mathcal{I}}^{(2)}},\widehat{f}_{n}\right]\leq C_{3}\frac{\log(n)}{nh}+2C^{2}_{1}h^{2s}_{n}.$
 By choosing the bandwidth as
-$h_{n}=\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+d}},$
+$h_{n}=\alpha\left(\frac{\log(n)}{n}\right)^{\frac{1}{2s+1}},$
 we complete the proof. ∎